# MathVC: An LLM-Simulated Multi-Character Virtual Classroom for Mathematics Education

### v1

@@ Abstract
Mathematical modeling (MM) is considered a fundamental skill for students in STEM disciplines. Practicing the MM skill is often the most effective when students can engage in group discussion and collaborative problem-solving. However, due to unevenly distributed teachers and educational resources needed to monitor such group activities, students do not always receive equal opportunities for this practice. Excitingly, large language models (LLMs) have recently demonstrated strong capability in both modeling mathematical problems (Wei et al., 2022) and simulating characters with different traits and properties (Park et al., 2023). Drawing inspiration from the advancement of LLMs, in this work, we present MathVC, the very first LLM-powered virtual classroom containing multiple LLM-simulated student characters, with whom a human student can practice their MM skill. To encourage each LLM character’s behaviors to be aligned with their specified math-relevant properties (termed “characteristics alignment”) and the overall conversational procedure to be close to an authentic student MM discussion (termed “conversational procedural alignment”), we proposed three innovations: integrating MM domain knowledge into the simulation, defining symbolic schema as the ground for character simulation, and designing a meta planner at the platform level to drive the conversational procedure. Through experiments and ablation studies, we confirmed the effectiveness of our simulation approach and showed the promise for MathVC to benefit real-life students in the future. [^1] Source code and examples will be available at our project website soon.
@@ 1 Introduction
Mathematical modeling (MM), the process of creating a mathematical representation of a real-world problem, is a critical skill for students pursuing STEM fields. Practicing this skill often requires students to engage in effective discussions of relevant concepts and collaboratively solve mathematics problems with their classmates (Vygotsky & Cole, 1978; Garfunkel & Montgomery, 2019). However, orchestrating such discussions and collaborative learning activities can be very challenging for teachers (Bush et al., 2020). As a consequence, the Opportunity Myth report (TNTP, 2018) has highlighted that students from marginalized communities where only limited teachers and educational resources are available, are less likely to have opportunities to participate in effective practices of MM skills.
@@
The rapidly developing large language models (LLMs) have shown huge potential to reshape the future of education (including Mathematics) (Denny et al., 2024). This potential stems from two recent advancements uniquely for LLMs. Firstly, LLMs present strong math reasoning capabilities in understanding and solving mathematics problems (Wei et al., 2022; Lu et al., 2023; Ahn et al., 2024). In particular, on the GSM8k dataset (Cobbe et al., 2021) representing middle-school mathematics level, the state-of-the-art GPT-4 model has achieved close to perfect performance (OpenAI et al., 2024). Secondly, by generating human-like thoughts and actions in explicit texts, autonomous LLM-based generative agents have showcased the capacity to simulate human activities (Park et al., 2023; Wang et al., 2023c; Zhou et al., 2023). Leveraging this capability, studies have explored simulating human behaviors in different scenarios, such as games (Wang et al., 2023a), social activities (Li et al., 2023b; Cheng & Chin, 2024), software engineering (Hong et al., 2023; Qian et al., 2023).
@@
Figure 1: An example demonstrating the application of MathVC, where students are presented a math problem (optionally with accompanied data) and engage in effective discussions on math modeling. Alice, Bob, and Charlie [^3] We use the three names as “placeholders” following the conversion in science and engineering literature (Wikipedia, n.d.); however, we envision the system to be used by various communities. are LLM-simulated student characters. A human student can choose to respond or skip the current turn.
@@
Observing the pressing need to enhance MM education for resource-limited communities and getting inspired by the striking progress made by LLMs, in this work, we propose MathVC, a Mathematics Virtual Classroom (Figure 1) featuring multiple LLM-simulated student characters, with whom a human student can discuss and collaboratively solve MM problems, which thus allows them to practice their MM skills without needing the teachers monitoring it. In particular, the platform focused on middle-school MM problems, a difficulty level that has been well handled by GPT-4. To the best of our knowledge, our research is the first work exploring the potential of leveraging LLMs and LLM-powered generative agents for mathematics education. We envision that such a platform can eventually be deployed as part of the middle-school curriculum, enabling effective MM learning to go beyond the space constraint in traditional classroom studies while enhancing students’ MM capability.
@@
We identified two unique challenges in the development of MathVC. The first challenge is characteristic alignment, i.e., aligning an LLM’s character simulation to the authentic characteristics of real human students. Middle-school students in real life could struggle with rapid and accurate modeling. They may initially produce incomplete or erroneous mathematical models and refine them iteratively as the discussion progresses. This process of gradually establishing mutual understanding with teammates, interleaved by continual self-reflection, is the key for students to improve their MM skills. However, in our preliminary experiments using GPT-4, a vanilla character simulation via prompt engineering the profile instruction to LLM cannot yield such an authentic and fine-grained student simulation. Instead, GPT-4 often directly gives a close-to-perfect solution, even when prompted to simulate students with poor math skills. Simulating how a student’s problem understanding and math modeling evolve during the collaboration, is even more difficult. The second challenge is conversational procedural alignment, i.e., aligning the overall conversational procedure to an authentic MM discussion among middle-school students. While the first challenge concerns individual characters’ simulation, this second challenge considers inter-character interaction. In practice, a collaborative problem-solving process typically consists of multiple stages, including problem understanding, task division, solution planning, plan execution, etc. (PISA, 2018). However, when we directly exposed the group of LLM characters to interact with each other, these LLM characters often bypassed the earlier stages and instead, initiated the conversation by immediately describing a modeling solution. For both of the two challenges, we conjecture that the vanilla simulation fails because LLMs were pre-trained in a question-answering style (e.g., they tend to directly give answers to the input mathematics questions), while lacking sufficient instructiblity to follow the prompt instruction and simulate the deliberation of a middle-school student. Despite existing exploration of LLM character simulations, we have not seen any prior work carefully exploring these two alignment challenges and the solutions to them.
@@
To tackle these challenges, we propose to integrate relevant theories of MM and collaborative problem-solving (CPS) (Standards, 2010; PISA, 2018) into the development of MathVC. Specifically, to enhance characteristics alignment, we propose using a symbolic character schema to represent the dynamic problem understanding process of a character. The character schema centers around “variables” that are essential for solving the mathematics problem, which are considered fundamental elements in MM theory (Standards, 2010). By considering it as a “ground”, we can simulate characters with more intricate behaviors, including how they make a mistake initially and then gradually fix it during the collaboration with other students. To enhance conversational procedural alignment, we further propose a meta planner at the platform level, which monitors the overall conversation and plans its procedure following the stages defined in the CPS theory (PISA, 2018).
@@ 2.1 Overview
Our goal is to develop a “virtual classroom” platform with multiple LLM-simulated student characters discussing and collaboratively solving MM tasks. We propose to design MathVC with a modular architecture as shown in Figure 3. Specifically, the system consists of two parts, i.e., meta planner for organizing the overall conversation and facilitating a smooth multi-stage student discussion, and character simulation for creating and updating individual student characters. All modules in MathVC are implemented by prompting an LLM (e.g., GPT-4) with specialized scripts, mostly in zero-shot (except for Dialogue Act Generator in one-shot). A complete workflow of the system is presented in Appendix A.
@@ 2.2 Symbolic Representations for Task and Character Modeling
A key innovation of MathVC lies in the use of symbolic schemas describing how characters form and gradually update their understanding of the mathematics problem. Specifically, we design two schema representations. Task Schema is a template describing elements that are necessary for solving a given MM task. As shown in Figure 2 (left), a task schema specifies how the original mathematics problem can be decomposed into multiple sub-tasks, and for each sub-task, what variables should be defined to resolve it. In other words, a task schema shows the ground-truth understanding and modeling of a mathematics problem. We define the schema centering around “variables” because they are considered the fundamental elements in MM (Standards, 2010). In practice, this schema can be automatically generated by prompting an LLM with the mathematics question and the ground-truth step-by-step solution following the Chain-of-Thought (Wei et al., 2022, CoT) formulation. Note that while a task schema could be challenging and time-consuming to manually annotate for teachers using this platform, the step-by-step solution is easy to compose or can be scraped from the answer notes of textbook exercises or directly generated by GPT-4.
Built on top of the task schema, we further devise Character Schema, a symbolic representation showing a student character’s understanding and modeling plan on the given mathematics problem. As shown in Figure 2 (right), a character schema is a task schema injected with errors that could be made by real human students. We consider a wide range of errors including misunderstanding on a sub-task or even omitting the sub-task, as well as any variable mistakes resulting from miscounting, incorrect modeling, or calculation errors. In mathematics education, materials such as the lecture notes in the Mathematics Assessment Project (MAP) (MAP, n.d.), often summarize common mistakes in each math problem, which can be directly used for simulating how students had incorrect solutions in the beginning. In MathVC, the character schema is synthesized by prompting an LLM to edit a task schema based on the student’s characteristics as well as a set of possible mistakes. Throughout the conversation with other students, a character will gradually update its character schema to reflect its dynamic thought process, as we will introduce in Section 2.4.
@@ 2.3 Meta Planning
The main goal of the meta-planning module is to perform meta-level control of the entire conversation to encourage conversational procedural alignment. It consists of three components: (1) Task Schema Generator: As Section 2.2 mentioned, the initial step of MathVC is to generate the task schema, which can then be edited to form character schemas. We instruct the LLM to produce the task schema with the specified format shown in Figure 2. A rule-based text checker will be applied to verify if the generated schema follows the format; if not, we regenerate a new one. (2) Collaboration Stage Monitor: Prior work (PISA, 2018) has observed that the collaborative problem-solving (CPS) process among students often consists of a sequence of stages, such as first establishing a shared task understanding and then establishing a team organization, as we summarize in Figure 3 and explained in Appendix B. To ensure that our simulated discussions closely resemble real CPS discussions, every dialogue will start from Stage 1, with the collaboration stage monitor dynamically determining whether the current stage is completed and whether the discussion should proceed to the next stage. This decision is made based on the dialogue history as well as a definition of every stage. When all stages are completed, we end the session. We note that simulating this multi-stage process is also important for human students to engage in effective MM practices, as it enables more scaffolded human and character interaction. (3) Dialogue Speaker Control: To facilitate smoother conversation interactions, we have developed a dialogue speaker control module that predicts the next speaker based on the dialogue context. For instance, if the last response is “Hey Bob, you made a mistake…”, the next speaker is likely Bob. The LLM captures this intuition by learning from the dialogue history, and it is prompted to output a single character name as the next speaker.
@@ 2.4 Character Simulation
Our simulation functions with four components (Figure 3, bottom). The first two components are responsible for generating and maintaining the character schema. (1) Character Schema Generator: As we introduced in Section 2.2, every simulated student character has its own character schema to reflect their thought. In our simulation, a character schema generator decides the specific mistake(s) a student character will make based on their traits and then edits the task schema into their initial character schema (Figure 2). To facilitate this editing, as shown in Appendix C, we represent the task schema as a Python dictionary and express every edit as a line of Python code that specifies the replacement of a variable or a sub-task description. (2) Character Schema Modifier: During real-life student discussions, a human student’s thought for a given question often evolves as the conversation progresses. We simulate these dynamics by continuously modifying the character schema based on the dialogue history. We prompt the LLM to decide whether the current schema needs to be modified. When it does, the modification will be achieved by generating a Python code that edits the previous character schema, similar to the initial character schema generation.
@@
Generating character responses is not a trivial task in our application. A major challenge we observed is their lack of diversity. For example, when we directly prompted LLMs to generate responses simulating how students discuss their modeling plans (i.e., Stage 3), all responses were uniformly assertive statements of a plan, without other forms of communication that are also commonly seen in authentic student discussions (e.g., question-answering). To address the issue, we propose a two-step response generation with the following two components. (3) Dialogue Act Generator: Before the response generation, we first prompt the LLM character to pick a “dialogue act” indicating the intent of their response. Dialogue acts are widely used in traditional dialogue systems (Stolcke et al., 2000; Li et al., 2016; Zhao et al., 2017). In our application, we define a set of dialogue acts for each conversational stage, such as “present a task understanding” and “ask for a clarifying question about a task understanding” for Stage 1, as shown in Figure 3; the complete definitions can be found in Appendix B. A dialogue act generator is then tasked to generate a dialogue act (e.g., “ask Bob a clarifying question about…”) based on the character schema, the dialogue history, and the set of dialogue acts for the current stage. We found that the use of dialogue acts could dramatically increase the response diversity when an LLM is explicitly prompted to reason about the proper response (e.g., asking a clarifying question when the character has a different modeling plan than their teammates, or generating a narrative answer when a character is asked for clarification). (4) Character Response Generator: Once deciding the dialogue act, a character response can be generated conditionally. However, even when the character schema is provided in the prompt, we still observe inconsistency between the character schema and the generated response (e.g., characters referring to inconsistent variable values than those in their schemas). To make the response follow the character schema better, instead of directly generating, we prompt an LLM to first retrieve relevant variables from the character schema, and then generate a response conditionally on the variables and their values. In practice, we found this prompt design to be very effective. Additionally, in the prompt, we instruct the LLM to mimic the language style of middle-school students (e.g., keeping the response short and colloquial). We present the prompts for the dialogue act generator and the response generator in Appendix I.
3 Experiment
3.1 Experimental Setup
Dataset and Experimental Setting We conducted experiments using middle-school mathematics problems from the MAP project (MAP, n.d.) and GSM8k (Cobbe et al., 2021). Questions from the MAP include both answers and common student mistakes, while those from the GSM8k only have answers; therefore, we manually annotate mistakes for them.
The most ideal evaluation of MathVC is to include real middle-school students in the loop, but this is very challenging and unsafe to implement (see our Ethics Statement). As our purpose of this work is mainly to showcase the feasibility of such a virtual classroom, during evaluation, we exclude real-student interaction and only focus on whether the multiple (3 in our experiments) simulated LLM characters can reasonably collaborate on solving MM tasks. The evaluation is then conducted by sampling multiple simulated dialogues and manually assessing their quality. Due to the cost of manual evaluation, we sample a small set of questions (5 as the dev set and 10 as the test set). For each question on the test set, we consider two student character configurations with varying MM skill combinations, and sample one multi-character dialogue in each for every system in comparison. We used GPT-4-turbo as the backbone LLM. We include more experimental details in Appendix D.
Baseline Systems We compare MathVC with four baselines. (1) Vanilla Simulation follows the prior work (Park et al., 2023) in prompt engineering the profile instruction to an LLM. Specifically, we directly present the LLM the math question, the student characteristics, and the dialogue history, and then instruct the LLM to generate responses as a middle-school student. (2) Domain-Specified Simulation further augments the vanilla simulation with domain-specific knowledge in its prompt, including the stages and possible acts in each stage, common student mistakes, the language style of a middle-school student, and so on. To demonstrate the effectiveness of our character schema and the meta planner, we include two more variants on top of domain-specified simulation, i.e., (2.1) w/ only character schema, where we include our proposed character schema in the prompt to the LLM and allow for continual schema modification, and (2.2) w/ only meta planner, where we augment the domain-specified simulation with our meta planner. The planned stage is added to the prompt for character response generation.
Evaluation Metrics The human evaluation was conducted by two student annotators. The annotators were instructed to rate each sampled dialogue on the degrees of characteristic alignment and conversational procedural alignment it reflected (1: worst; 4: best). We report an average score. More details and the annotation guideline are presented in Appendix H.

### v2

@@ Abstract
Mathematical modeling (MM) is considered a fundamental skill for students in STEM disciplines. Effective MM practice often involves group discussions and collaborative problem-solving. However, due to unevenly distributed teacher resources and various circumstances, students do not always receive equally effective opportunities for such engagement. Excitingly, large language models (LLMs) have recently demonstrated strong capability in both modeling mathematical problems and simulating characters with different traits and properties. Drawing inspiration from the advancement of LLMs, in this work, we present MathVC, the very first LLM-powered virtual classroom containing multiple LLM-simulated student characters, with whom a human student can practice their MM skill. To encourage each LLM character’s behaviors to be aligned with their specified math-relevant properties (termed “characteristics alignment”) and the overall conversational procedure to be close to an authentic student MM discussion (termed “conversational procedural alignment”), we proposed three innovations: integrating MM domain knowledge into the simulation, defining symbolic schema as the ground for character simulation, and designing a meta planner to guide the conversational process. Our experiments and ablation studies validate our simulation method and demonstrate MathVC’s potential to enhance real-world student learning.
@@ 1 Introduction
Mathematical modeling (MM), the process of creating a mathematical representation of a real-world problem, is a critical skill for students pursuing STEM fields. Practicing this skill often requires students to engage in effective discussions of relevant concepts and collaboratively solve mathematics problems with their classmates (Vygotsky and Cole 1978; Garfunkel and Montgomery 2019). However, due to the limited amount of teachers who are well-trained to facilitate collaborative MM discussions, students from marginalized communities do not often receive equal opportunities to effectively practice their MM skills, as reported by the Opportunity Myth report (TNTP 2018). On the other hand, certain student populations, such as students suffering from mental barriers (e.g., anxiety) in peer interaction, or students coming from populations that historically show inferior performance in STEM disciplines and thus display low self-esteem, may confront further challenges that prevent them from effectively engaging in collaborative MM discussions.
@@
The rapidly developing large language models (LLMs) have shown huge potential to reshape the future of education (including Mathematics) (Denny et al. 2024). This potential stems from two recent advancements unique to LLMs. Firstly, LLMs present strong math reasoning capabilities in understanding and solving mathematics problems (Wei et al. 2022; Lu et al. 2023; Ahn et al. 2024). In particular, on the GSM8k dataset (Cobbe et al. 2021) representing middle-school mathematics level, the state-of-the-art GPT-4 model has achieved close to perfect performance (OpenAI et al. 2024). Secondly, by generating human-like thoughts and actions in explicit texts, autonomous LLM-based generative agents have showcased the capacity to simulate human activities (Park et al. 2023; Wang, Chiu, and Chiu 2023; Zhou et al. 2023). Leveraging this capability, studies have explored simulating human behaviors in different scenarios, such as games (Wang et al. 2023a), social activities (Li, Zhang, and Sun 2023; Cheng and Chin 2024), and software engineering (Hong et al. 2023; Qian et al. 2023).
@@
Figure 1: The current interface of MathVC, where students are presented a math problem and engage in effective discussions on math modeling. A human student can choose to respond or skip the current turn. The virtual peers, Alice, Bob, and Charlie, are all LLM-simulated characters.
00footnotetext: We use the three names as “placeholders” following the conversion in science and engineering literature (Wikipedia n.d.); however, we envision the system to be used by various communities.
@@
Observing the pressing need to enhance MM education for resource-limited and inherently disadvantaged communities and getting inspired by the striking progress made by LLMs, in this work, we propose MathVC, a Mathematics Virtual Classroom (Figure 1) featuring multiple LLM-simulated student characters, with whom a human student can discuss and collaboratively solve MM problems. In particular, the platform focused on middle-school math problems, a difficulty level that current advanced LLMs have handled well. Such a platform thus allows students to practice their MM skills without needing the teachers to monitor the collaborative discussions. Students having difficulty in real-life peer interaction can also receive effective practices through interacting with virtual peers. To the best of our knowledge, our research is the first work exploring the potential of leveraging multi-LLM agents for mathematics education. Notably, MathVC allows teachers to flexibly configure the attributes (e.g., backgrounds and math skill levels) of the virtual students based on the human student’s performance, such that the human student can improve their MM skills from the interaction. We consider this configuration as a new paradigm of teacher supervision at the meta-level, as opposed to directly supervising the student discussion. We envision that MathVC can eventually be deployed as part of the middle-school curriculum, enabling effective MM learning to go beyond the space constraint in traditional classroom studies while enhancing students’ MM capability.
@@
To better simulate a real discussion, we identified two unique challenges in the development of MathVC. The first challenge is characteristic alignment, i.e., aligning an LLM’s character simulation to the authentic characteristics of real human students. Middle-school students in real life could struggle with rapid and accurate modeling. They may initially produce incomplete or erroneous mathematical models and refine them iteratively as the discussion progresses. This process of gradually establishing mutual understanding with teammates, interleaved by continual self-reflection, is the key for students to improve their MM skills. However, in our preliminary experiments using LLMs, a vanilla character simulation via prompt engineering of the profile instruction to the LLM cannot yield such an authentic and fine-grained student simulation. Instead, it often directly gives a close-to-perfect solution, even when prompted to simulate students with poor math skills. Simulating how a student’s problem understanding and modeling evolve during the collaboration, is even more difficult. On the other hand, we also aim to simulate students at different math skill levels. One study found that students often benefit more from interacting with mixed-ability peers rather than those with significantly higher achievement (Lou 2013). Simulating students at different performance levels, especially low-achieving students, encourages the student users to establish a more robust concept understanding and participate more actively in discussions, such as by correcting the mistakes made by the simulated peers. We consider this simulation a goal of characteristic alignment as well.
The second challenge is conversational procedural alignment, i.e., aligning the overall conversational procedure to an authentic MM discussion among middle-school students. While the first challenge concerns individual characters’ simulation, this second challenge considers inter-character interaction. In practice, a collaborative problem-solving process typically consists of multiple stages, including problem understanding, task division, solution planning, plan execution, etc. (PISA 2018). There is evidence suggesting that structured collaborative learning, where students work together with multiple discussion stages, clear roles, and shared goals, tends to be more effective than simply having the most knowledgeable student (in our case, this is a vanilla LLM simulated student) give direct answers to the group (Rajaram and Rajaram 2021). It can increase student engagement, stimulate positive emotions and excitement, and have a positive impact on student academic performance. However, when we directly exposed the group of LLM characters to interact with each other, these LLM characters often bypassed the earlier stages and instead, initiated the conversation by immediately describing a solution.
@@
To tackle these challenges, we propose to integrate relevant theories of MM and collaborative problem-solving (CPS) (Standards 2010; PISA 2018) into the development of MathVC. Specifically, to enhance characteristics alignment, we propose using a symbolic character schema to represent the dynamic problem understanding process of a character. The character schema centers around “variables” that are essential for solving the mathematics problem, which are considered fundamental elements in MM theory (Standards 2010). By considering it as a “ground”, we can simulate characters with more intricate behaviors, including how they make a mistake initially and then gradually fix it during the collaboration with other students. Unlike us, previous LLM-based simulations have mostly used unstructured texts to represent a character’s memory (Park et al. 2023). However, our preliminary experiment showed that the unstructured text representation often yielded hallucinations in simulating a student’s math thinking (e.g., a sharp change in a student’s problem understanding without contexts) and also caused higher latency. Using symbolic schemas can fundamentally solve these issues and enhance “characteristics alignment”. To enhance conversational procedural alignment, we further propose a meta planner at the platform level, which monitors the overall conversation and plans its procedure following the stages defined in the CPS theory (PISA 2018).
@@ 2.1 Overview
Our goal is to develop a “virtual classroom” platform with multiple LLM-simulated student characters discussing and collaboratively solving MM tasks. We propose to design MathVC with a modular architecture as shown in Figure 3. Specifically, the system consists of two parts, i.e., meta planner for organizing the overall conversation and facilitating a smooth multi-stage student discussion, and character simulation for creating and updating individual student characters. All modules in MathVC are implemented by prompting an LLM (e.g., GPT-4) with specialized scripts, mostly in zero-shot (except for Dialogue Act Generator in few-shot).
@@ 2.2 Symbolic Representations for Task and Character Modeling
A key innovation of MathVC lies in the use of symbolic schemas describing how characters form and gradually update their understanding of the mathematics problem. Specifically, we design two schema representations. Task Schema is a template describing elements that are necessary for solving a given MM task. As shown in Figure 2 (left), a task schema specifies how the original mathematics problem can be decomposed into multiple sub-tasks, and for each sub-task, what variables should be defined to resolve it. In other words, a task schema shows the ground-truth understanding and modeling of a mathematics problem. We define the schema centering around “variables” because they are considered the fundamental elements in MM (Standards 2010). In practice, this schema can be automatically generated by prompting an LLM with the mathematics question and the ground-truth step-by-step solution following the Chain-of-Thought (Wei et al. 2022, CoT) formulation. Note that while a task schema could be challenging and time-consuming to manually annotate for teachers using this platform, the step-by-step solution is easy to compose or can be scraped from the answer notes of textbook exercises or directly generated by GPT-4.
Built on top of the task schema, we further devise Character Schema, a symbolic representation showing a student character’s understanding and modeling plan on the given mathematics problem. As shown in Figure 2 (right), a character schema is a task schema injected with errors that could be made by real human students. We consider a wide range of errors including misunderstanding on a sub-task or even omitting the sub-task, as well as any variable mistakes resulting from miscounting, incorrect modeling, or calculation errors. In mathematics education, materials such as the lecture notes in the Mathematics Assessment Project (MAP) (MAP n.d.), often summarize common mistakes in each math problem, which can be directly used for simulating how students had incorrect solutions in the beginning. In MathVC, the character schema is synthesized by prompting an LLM to edit a task schema based on the student’s characteristics as well as a set of possible mistakes. Throughout the conversation with other students, a character will gradually update its character schema to reflect its dynamic thought process, as we will introduce in Section 2.4.
@@ 2.3 Meta Planning
The main goal of the meta-planning module is to perform meta-level control of the entire conversation to encourage conversational procedural alignment. It consists of three components: (1) Task Schema Generator: As Section 2.2 mentioned, the initial step of MathVC is to generate the task schema, which can then be edited to form character schemas. We instruct the LLM to produce the task schema with the specified format shown in Figure 2. A rule-based text checker will be applied to verify if the generated schema follows the format; if not, we regenerate a new one. (2) Collaboration Stage Monitor: Prior work (PISA 2018) has observed that the collaborative problem-solving (CPS) process among students often consists of a sequence of stages, such as first establishing a shared task understanding and then establishing a team organization, as we summarize in Appendix A. To ensure that our simulated discussions closely resemble real CPS discussions, every dialogue will start from Stage 1, with the collaboration stage monitor dynamically determining whether the current stage is completed and whether the discussion should proceed to the next stage. This decision is made based on the dialogue history as well as a definition of every stage. When all stages are completed, we end the session. We note that simulating this multi-stage process is also important for human students to engage in effective MM practices, as it enables more scaffolded human and character interaction. (3) Dialogue Speaker Control: To facilitate smoother conversation interactions, we have developed a dialogue speaker control module that predicts the next speaker based on the dialogue context. For instance, if the last response is “Hey Bob, you made a mistake…”, the next speaker is likely Bob. The LLM captures this intuition by learning from the dialogue history, and it is prompted to output a single character name as the next speaker.
Human-System Interaction Design
Our current system interacts with the user through the dynamic control with the Dialogue Speaker Control module. If the next speaker candidates include only the user, the system will wait indefinitely for the human student’s response. If the next speaker candidates include both agents and the user, the system will wait for 10 seconds; if the student does not input any characters during this time, it will be considered the agent’s turn, and the corresponding agent’s response will be output. If the student provides input, the conversation will proceed based on the student’s input. If the next speaker candidates include only agents, the system will directly output the agent’s response.
@@ 2.4 Character Simulation
Our simulation functions with four components (Figure 3, bottom). The first two components are responsible for generating and maintaining the character schema. (1) Character Schema Generator: As we introduced in Section 2.2, every simulated student character has its own character schema to reflect their thought. In our simulation, a character schema generator decides the specific mistake(s) a student character will make based on their traits and then edits the task schema into their initial character schema (Figure 2). To facilitate this editing, we represent the task schema as a Python dictionary and express every edit as a line of Python code that specifies the replacement of a variable or a sub-task description. (2) Character Schema Modifier: During real-life student discussions, a human student’s thought for a given question often evolves as the conversation progresses. We simulate these dynamics by continuously modifying the character schema based on the dialogue history. We prompt the LLM to decide whether the current schema needs to be modified. When it does, the modification will be achieved by generating a Python code that edits the previous character schema, similar to the initial character schema generation.
@@
Generating character responses is not a trivial task in our application. A major challenge we observed is their lack of diversity. For example, when we directly prompted LLMs to generate responses simulating how students discuss their modeling plans (i.e., Stage 3), all responses were uniformly assertive statements of a plan, without other forms of communication that are also commonly seen in authentic student discussions (e.g., question-answering). To address the issue, we propose a two-step response generation with the following two components. (3) Dialogue Act Generator: Before the response generation, we first prompt the LLM character to pick a “dialogue act” indicating the intent of their response. Dialogue acts are widely used in traditional dialogue systems (Stolcke et al. 2000; Li et al. 2016; Zhao, Zhao, and Eskenazi 2017). In our application, we define a set of dialogue acts for each conversational stage, such as “present a task understanding” and “ask for a clarifying question about a task understanding” for Stage 1, as shown in Figure 3; the complete definitions can be found in Appendix A. A dialogue act generator is then tasked to generate a dialogue act (e.g., “ask Bob a clarifying question about…”) based on the character schema, the dialogue history, and the set of dialogue acts for the current stage. We found that the use of dialogue acts could dramatically increase the response diversity when an LLM is explicitly prompted to reason about the proper response (e.g., asking a clarifying question when the character has a different modeling plan than their teammates, or generating a narrative answer when a character is asked for clarification). (4) Character Response Generator: Once deciding the dialogue act, a character response can be generated conditionally. However, even when the character schema is provided in the prompt, we still observe inconsistency between the character schema and the generated response (e.g., characters referring to inconsistent variable values than those in their schemas). To make the response follow the character schema better, instead of directly generating, we prompt an LLM to first retrieve relevant variables from the character schema, and then generate a response conditionally on the variables and their values. In practice, we found this prompt design to be very effective. Additionally, in the prompt, we instruct the LLM to mimic the language style of middle-school students (e.g., keeping the response short and colloquial). We present the prompts for the dialogue act generator and the response generator in Appendix I.
3 System Evaluation